salopiantree's posterous

salopiantree's posterous

salopiantree  //  A site purely for testing for other sites

Mar 17 / 4:57pm

PhPList.com Bulletin 1

ICT Daily Bulletin 16th March 2011

Good morning all. Please find below a list of the recent issues facing ICT. If you have anything you think should be announced to our customers or the rest of ICT, please send them to incident.manager@imperial.ac.uk

Our Service Status page (which shows you any current major issues relating to ICT services) now has a RSS feed associated to it. Please go to http://www.imperial.ac.uk/ict/servicedesk/status/ for more information. Any feedback you wish to provide concerning this page can be emailed to incident.manager@imperial.ac.uk.

 

Today:

Incident-562241:  SharePoint was unavailable overnight.  The cause is being investigated.

16:30 - 18:30

E-Learning (Blackboard, Echo360) at risk.

17:00 - 18:00

CHG-561822:  Upgrade SAN CISCO MDS switches to 5.0.4, no downtime.  Subject to approval.

18:00 - 23:00

IC Analytics at risk, during which:

18:00 - 18:30

CHG-559637:  Minor: BIPA w11 - two additional fields in BIEE purchasing, downtime to ICA.

21:00 - 21:05

CHG-561527:  Database switchovers on Exchange 2010, no downtime to user expected.

 

Tomorrow morning:

No planned works notified to the ICT Service Desk.

Other notes:

14th March, 09:00 - 18th March, 17:00

CHG-561025:  Install rollup 3 for Exchange server 2010 service pack 1, no downtime.

14th March - 27th March

Spring SOLE suite of surveys open for all undergraduate and post graduates.

TRAFFIC LIGHT REPORT

Key Service

Previous 24 Hrs

Live Status

Authentication and Directory Services

OK

OK

Data Centre

OK

OK

Email

OK

OK

File / Backup

OK

OK

A Key Business Application (ICIS, OSS, InfoEd, etc)

OK

OK

Library

OK

OK

Network Infrastructure

OK

OK

Printing

OK

OK

Security Services

OK

OK

Student Halls Network

OK

OK

Teaching and Public Access Library Clusters

OK

OK

Teaching Services

OK

OK

Telephony

OK

OK

Web Services

SharePoint

OK

Other issues of note

OK

Red Test

NEW AND SIGNIFICANT MARVAL PROBLEMS

The list of outstanding problem requests is available via the online Service Status page. Please visit the site to see the list of Problem Records. This provides a more in-depth picture of open Problems ICT is currently working on than can be displayed in our bulletin. You can also subscribe to the Problem RSS feed.

FUTURE

Friday 18th March, 18:00 until Sunday 20th March, 18:00

ICIS unavailable

Friday 25th March, 13:00 until Monday 28th March, 07:00

Electrical power shutdown to level 3, East corridor, Skempton building.  Computing room 317 and data network service will not be affected.

Saturday 26th March, 07:30 until Sunday 27th March, 20:00

RCS1 will be closed and will have no power.  All sockets with prefixes of 166,216 and 279 will have no data services.

Saturday 2nd April, 21:00 - 00:00

Power outage to Growth Lab, 105, RCS1.

HISTORY

Tuesday 15th March:

07:00 - 09:00

Network at risk, during which:

TSK-561317:  Enable IPv6 connectivity on Brompton hospital and Emmanuel Kay Building, no downtime

09:00 - 13:00

11:30

CHG-550643:  Change group type from global to universal for ICT staff and groups in active directory - administrative department and medicine, no downtime

17:00 - 21:00

InfoEd and Discoverer for small systems at risk

18:00 - 21:00

Apex at risk, during which:

18:00 - 18:30

CHG-559370:  APEX temp bank, 2011-W11.  Amendment to temp bank for non-SoV staff, downtime to temp bank.

21:00

CHG-561527:  Database switchovers on Exchange 2010

Deveena Raithatha

ICT Service Centre

 

ICT Daily Bulletin 16th March 2011

Good morning all. Please find below a list of the recent issues facing ICT. If you have anything you think should be announced to our customers or the rest of ICT, please send them to incident.manager@imperial.ac.uk

Our Service Status page (which shows you any current major issues relating to ICT services) now has a RSS feed associated to it. Please go to http://www.imperial.ac.uk/ict/servicedesk/status/ for more information. Any feedback you wish to provide concerning this page can be emailed to incident.manager@imperial.ac.uk.

Today:

Incident-562241:  SharePoint was unavailable overnight.  The cause is being investigated.

16:30 - 18:30

E-Learning (Blackboard, Echo360) at risk.

17:00 - 18:00

CHG-561822:  Upgrade SAN CISCO MDS switches to 5.0.4, no downtime.  Subject to approval.

18:00 - 23:00

IC Analytics at risk, during which:

18:00 - 18:30

CHG-559637:  Minor: BIPA w11 - two additional fields in BIEE purchasing, downtime to ICA.

21:00 - 21:05

CHG-561527:  Database switchovers on Exchange 2010, no downtime to user expected.

Tomorrow morning:

No planned works notified to the ICT Service Desk.

Other notes:

14th March, 09:00 - 18th March, 17:00

CHG-561025:  Install rollup 3 for Exchange server 2010 service pack 1, no downtime.

14th March - 27th March

Spring SOLE suite of surveys open for all undergraduate and post graduates.

TRAFFIC LIGHT REPORT

Key Service

Previous 24 Hrs

Live Status

Authentication and Directory Services

OK

OK

Data Centre

OK

OK

Email

OK

OK

File / Backup

OK

OK

A Key Business Application (ICIS, OSS, InfoEd, etc)

OK

OK

Library

OK

OK

Network Infrastructure

OK

OK

Printing

OK

OK

Security Services

OK

OK

Student Halls Network

OK

OK

Teaching and Public Access Library Clusters

OK

OK

Teaching Services

OK

OK

Telephony

OK

OK

Web Services

SharePoint

OK

Other issues of note

OK

Red Test

NEW AND SIGNIFICANT MARVAL PROBLEMS

The list of outstanding problem requests is available via the online Service Status page. Please visit the site to see the list of Problem Records. This provides a more in-depth picture of open Problems ICT is currently working on than can be displayed in our bulletin. You can also subscribe to the Problem RSS feed.

FUTURE

Friday 18th March, 18:00 until Sunday 20th March, 18:00

ICIS unavailable

Friday 25th March, 13:00 until Monday 28th March, 07:00

Electrical power shutdown to level 3, East corridor, Skempton building.  Computing room 317 and data network service will not be affected.

Saturday 26th March, 07:30 until Sunday 27th March, 20:00

RCS1 will be closed and will have no power.  All sockets with prefixes of 166,216 and 279 will have no data services.

Saturday 2nd April, 21:00 - 00:00

Power outage to Growth Lab, 105, RCS1.

HISTORY

Tuesday 15th March:

07:00 - 09:00

Network at risk, during which:

TSK-561317:  Enable IPv6 connectivity on Brompton hospital and Emmanuel Kay Building, no downtime

09:00 - 13:00

11:30

CHG-550643:  Change group type from global to universal for ICT staff and groups in active directory - administrative department and medicine, no downtime

17:00 - 21:00

InfoEd and Discoverer for small systems at risk

18:00 - 21:00

Apex at risk, during which:

18:00 - 18:30

CHG-559370:  APEX temp bank, 2011-W11.  Amendment to temp bank for non-SoV staff, downtime to temp bank.

21:00

CHG-561527:  Database switchovers on Exchange 2010

Deveena Raithatha

ICT Service Centre